OpenStudy (anonymous):

PLEASE HELP! Will Medal! Which is true? Light waves do not change speed as they travel from air through water. Light waves speed up as they travel from the air through the water Light waves slow down as they travel from the air through the water.

OpenStudy (anonymous):

@texaschic101 @dan815 @Luigi0210 @wio @sammixboo @nincompoop @uri @kropot72

OpenStudy (michele_laino):

the speed v of light in a medium whose refractive index is n, is given by the subsequent formula: \[\Large v = \frac{c}{n}\] where c is the light speed in vacuum Now the refractive index of air is about 1. Next the refractive index of water is greater than 1, so we can write: \[\Large {v_{WATER}} = \frac{c}{{{n_{WATER}}}}\] \[\Large {n_{WATER}} > 1\] what can you conclude?

rvc (rvc):

@Michele_Laino it will low dows

OpenStudy (michele_laino):

yes since: being \[\Large {n_{WATER}} > 1\] then: \[\Large {v_{WATER}} = \frac{c}{{{n_{WATER}}}} < c\]

rvc (rvc):

yay thanks :)

OpenStudy (michele_laino):

:)